Circular No. 7298

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 U4 (CATALINA-SKIFF) Yesterday T. B. Spahr, Lunar and Planetary Laboratory, reported the

discovery by the Catalina Sky Survey of a slow-moving object that was independently discovered today by B. A. Skiff (measurer B. W. Koehn) by the LONEOS survey. In response to Skiff’s alert, R. L. Millis and L. H. Wasserman, on a 5-min R-band exposure with the Perkins 1.8-m reflector, detected a coma extending 8′′ southeastward from the nucleus. After a posting in The NEO Confirmation Page, M. Tichý and Z. Moravec, Kleť, also reported that the object had an 8′′ coma. The discovery observations follow, with very preliminary parabolic elements and an ephemeris from MPEC 1999-V04.

1999 UT α_{2000} δ_{2000} m_{1} Observer
Oct. 31.24836 2^{h}08^{m}48.^{s}20 +20^{o}15′56′′.8 17.0 Spahr
Nov. 1.27779 2 08 14.92 +20 16 35.4 17.4 Skiff
T = 2001 Sept.14.404 TT  ω = 77.^{o}087
Ω = 31.540 }2000.0
q = 4.77674 AU  i = 47.304
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 29 2^{h}10.^{m}01 +20^{o}14.^{′}5 6.065 7.052 173.^{o}0 1.^{o}0 16.9
Nov. 3 2 07.32 +20 17.6 6.042 7.026 172.2 1.1 16.9
8 2 04.64 +20 20.2 6.027 7.000 168.4 1.6 16.9
13 2 02.01 +20 22.4 6.020 6.974 163.5 2.3 16.8
18 1 59.45 +20 24.3 6.020 6.949 158.3 3.0 16.8

SUPERNOVA 1999ej IN NGC 495 S. Jha, P. Challis, P. Garnavich, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum obtained by M. Calkins with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) on Oct. 30.27 UT shows SN 1999ej to be a normal type-Ia supernova near maximum light. Adopting the NED recession velocity for NGC 945 of 4414 km/s yields expansion velocities for Si II (rest 635.5 nm) and Ca II (H+K) of 10 500 km/s and 15 700 km/s, respectively. The estimated age of the supernova based on the observed spectral features (Riess et al. 1998, Ap.J. 504, 935) is 1.5 ± 3 days before maximum light.

1999 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green